Circular No. 6155

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

AL COMAE BERENICES J. A. Mattei, AAVSO, reports that this cataclysmic variable is in out-

burst for the first time since March 1975 (IAUC 2760). Available visual magnitude estimates: Apr. 3.804 UT, 15.7: (L. Szentaskó, Veresegyház, Hungary); 4.97, [13.7 (P. Schmeer, Bischmisheim, Germany); 5.149, 14.9 (D. York, Abiquiu, NM); 5.600, 12.7 (T. Cragg, Coonabarabran, New South Wales); 5.799, 12.4 (Szentaskó); 5.83, 12.2 (Schmeer).

GRAVITATIONAL MICROLENSING EVENTS The MACHO collaboration (A. Alcock, T. Axelrod, R. Allsman, D.

Bennett, K. Cook, K. Freeman, K. Griest, S. Marshall, B. Peterson, M. Pratt, P. Quinn, D. Reynolds, A. Rodgers, C. Stubbs, and W. Sutherland; Lawrence Livermore National Laboratory, Center for Particle Astrophysics, and Mount Stromlo Observatory) reports the discovery of likely gravitational microlensing events in progress toward the galactic bulge: “The stars in question and their positions are as follows: event 95-1, α = 17^{h}59^{m}00^{s}.6, δ = −28^{o}36′58′′ (equinox 2000.0); event 95-2, α = 18^{h}08^{m}25^{s}.2, δ = −27^{o}58′38′′. Both stars were constant at magnitudes R = 17.9 and V = 19.0 during 1993 and have brightened by 1.6 and 0.5 mag, respectively, as of 1995 Apr. 4. Microlensing fits to data obtained with the Mt. Stromlo 1.27-m telescope indicate that event 95-1 may reach a high amplitude between Apr. 10 and 12 or reach its peak at a more modest amplitude prior to Apr. 10. Event 95-2 appears to be a longer-timescale event that may not get brighter than its present brightness. Accurate photometry near the peak of event 95-1 may yield valuable information about the nature of the microlensing object. Finding charts and additional information regarding the MACHO Alert System are available by anonymous ftp, address

darkstar.astro.washington.edu, in the directory macho/Alert; this di-

rectory will be periodically updated with new information as new microlensing events are discovered. Potential observers are requested to contact macho@astro.washington.edu and/or D. Bennett at telephone 510-423-0656 (USA), to coordinate observations.”

NOVA AQUILAE 1995 Visual magnitude estimates (cf. IAUC 6146): Mar. 9.23 UT, 9.7 (A.

Pereira, Cabo da Roca, Portugal); 14.52, 10.4 (P. Collins, Scottsdale, AZ); 23.50, 10.3 (Collins); Apr. 3.18, 10.4 (Pereira).

1995 April 5 Daniel W. E. Green